# Neutrino Heating in 1D, 2D, and 3D core-collapse supernovae: characterizing the explosion of high-compactness stars

### v1

@@ 2.2 Neutrino heating in 1D+ models
To illustrate the effects of convection and neutrino heating in the 1D+ STIR model we show in Figure 1 the time and radial dependence of the rate of generation of turbulent energy (i.e. $\epsilon_{g}$) as well as the rate of net energy deposition due to neutrinos (i.e. $\rho\partial\epsilon_{\nu}/\partial t$) for 9 different simulations from Boccioli & Fragione (2024). We selected simulations from Boccioli (2024a) characterized by different values of pre-supernova compactnesses $\xi_{2.0}$. The properties of the selected progenitors are shown in Table 1.
@@
Table 1: Summary of the properties of the progenitors used in Section 4. The simulations were performed in Boccioli (2024a).
Progenitors with larger compactness are characterized by shallower density profiles, and therefore larger mass accretion rates. As a consequence, they generate larger neutrino heating, while at the same time, they also generate larger ram pressure onto the shock. In the remainder of this paper, we will show that high-compactness progenitors, in the presence of $\nu$-driven convection, are able to generate large neutrino heating very efficiently, which has a direct impact on explodability.
@@
which represent the turbulent generation and dissipation, respectively. Notice that, in principle, one expects $E_{g}\approx E_{k}$, since the overall turbulence generated should balance the turbulence dissipated, as shown by Murphy et al. (2013). This is however not necessarily true in STIR, and instead, $E_{g}$ is about $10\%$ larger than $E_{k}$.
@@ 3 Comparison between STIR and multi-dimensional simulations
To illustrate how the STIR model compares to multi-dimensional simulations, we compare neutrino heating in STIR with neutrino heating in state-of-the-art 2D and 3D Fornax and 2D FLASH simulations. Specifically, we will analyze 98 2D and 20 3D Fornax simulations and 55 2D FLASH simulations. The 2D Fornax simulations (Vartanyan & Burrows, 2023) were performed on solar metallicity progenitors from Sukhbold et al. (2016) with initial masses between 9 and 12 $M_{\odot}$, and on solar metallicity progenitors from Sukhbold et al. (2018) with initial masses between 12 and 26.99 $M_{\odot}$. The 3D Fornax simulations (Burrows et al., 2020) were performed on solar metallicity progenitors from Sukhbold et al. (2016), with masses ranging from 12 to 60 $M_{\odot}$, as well as a $25M_{\odot}$ progenitor from Sukhbold et al. (2018). All of the Fornax simulations adopted the SFHo equation of state. The 2D FLASH simulations were performed on different progenitors with masses from $12-60M_{\odot}$ from Sukhbold et al. (2016) as well as a few from Woosley & Heger (2007) and some with artificially reduced mass loss from Sukhbold et al. (2018). Of these progenitors, 37 were simulated using the SFHo EOS, whereas the rest were simulated using other EOSs from Schneider et al. (2019). The FLASH simulations are collected from various studies but also from publications in preparation (Eggenberger Andersen et al., 2021, 2024; Li, 2024; Andresen et al., 2024).
@@
Several groups have performed comparisons between different codes, although usually it has been done for simulations in either 1D (O’Connor et al., 2018), 2D (Pan et al., 2019), or 3D (Cabezón et al., 2018; Glas et al., 2019), and more rarely have 1D simulations been thoroughly compared to multi-dimensional ones (Just et al., 2018; Richers et al., 2017; Nagakura et al., 2020). The latter is a particularly difficult comparison to perform given that 1D simulations do not explode, whereas multi-dimensional ones often do, and therefore the most significant differences between 1D and multi-D results can be attributed to the different outcome rather than specific multi-dimensional effects.
@@
In this section, we focus on comparing neutrino heating in 1D, 1D+, 2D, and 3D simulations. We adopt simulations already performed with other studies in mind, and therefore neutrino transport, opacities, and overall simulation setups are not necessarily the same across codes, which prevents a detailed and thorough comparison study, which is however beyond the scope of this paper. Instead, we specifically focus on differences in neutrino heating, which is the subject of the present study.
We analyze simulations performed with three different codes: GR1D, Fornax, and FLASH. All of the simulations were run using the SFHo equation of state (Hempel & Schaffner-Bielich, 2010; Steiner et al., 2013), except for 18 2D FLASH simulations, run with different EOSs as described above. The neutrino transport and opacities are not necessarily the same among the different codes, since the simulations were not originally run with this comparison in mind. We refer to the original papers listed above for details regarding the specific neutrino treatments in the different codes. Therefore, some of the discrepancies between 1D+ and multi-D, as well as discrepancies between the 2D FLASH and Fornax simulations, are to be expected.
@@
In general, neutrino heating (but also several other quantities) in 2D simulations tends to be a lot noisier than in 3D, even though the neutrino luminosities of 2D and 3D Fornax simulations agree very well (Vartanyan & Burrows, 2023). Because of the imposed axial symmetry, 2D simulations are characterized by much larger structures than their 3D counterparts, and therefore neutrino heating does not average out as much compared to 3D.
@@
For the low-compactness progenitor, the 2D Fornax simulations produce a larger neutrino heating at early times, compared to both the 2D FLASH and 3D Fornax simulations, which instead seem to agree relatively well. This is seen for all of the lower-compactness progenitors, as can be inferred from the larger $\dot{Q}_{\rm max}$ in Figure 3 (see Section 3 for more details) for 2D Fornax models compared to 3D Fornax and 2D FLASH models. The larger neutrino heating in the 2D Fornax simulations can be caused by a number of factors, such as the different geometry and resolution. Minor differences in neutrino transport and different opacities can also cause some differences in neutrino heating, but a detailed analysis of the source of this discrepancy is beyond the scope of this work.
@@
An important detail worth mentioning is that, depending on how the gain region is defined, certain computational zones might be included or excluded from the integral in Eq. (8). For example, if one defines the gain region as the region with positive net neutrino heating and densities lower than $3\times 10^{10}\ {\rm g/cm^{3}}$, certain zones close to the PNS might be included in the integral, even though they do not influence the neutrino heating in the vicinity of the shock, which ultimately is what Eq. (8) should reflect. Therefore, we define the gain region in the same way for all of the simulations analyzed, i.e. the region where net neutrino heating is positive, density is below $3\times 10^{10}\ {\rm g/cm^{3}}$ and entropy per baryon is above $6$. The latter condition ensures that low-entropy material, i.e. material that has cooled via neutrinos and is settling onto the PNS, is not included in the integral. This simple condition can decrease the neutrino heating in the 2D simulations by up to $\sim 10\%$.
@@ 3.1 Compactness dependence of neutrino heating
In this section, we will focus on several suites of simulations in 1D, 1D+, 2D and 3D, to study how neutrino heating depends on the progenitor compactness. For that purpose, we will use 20, 98, and 55 simulations from the 3D Fornax, 2D Fornax, and 2D FLASH models described above, respectively. In addition to that, we selected several suites of both 1D and 1D+ simulations ran with GR1D: 309 FRANEC progenitors (Chieffi et al., 2021) from Boccioli et al. (2023); 341 KEPLER progenitors (Sukhbold et al., 2016; Woosley et al., 2002) from Boccioli & Fragione (2024); 88 (only 1D+ simulations) KEPLER progenitors (Sukhbold et al., 2018) simulated with GR1D, which are the same progenitors as the 2D Fornax simulations by Vartanyan & Burrows (2023). Finally, we also selected 148 1D FLASH simulations (Segerlund et al., 2021) of KEPLER progenitors from Sukhbold et al. (2016) with masses between 9 and 120 $M_{\odot}$, as well as 40 1D+ FLASH simulations (Couch et al., 2020) with masses from 9 to $30M_{\odot}$. In total, we therefore have 173 multi-D simulations, 778 1D+ simulations, and 798 1D simulations, for a total of $\sim 1500$ simulations.
@@
Figure 3 shows $\dot{Q}_{\nu}^{\rm max}$ as as a function of the pre-supernova compactness for the $\sim 1500$ simulations described above. The bottom panel is color-coded based on explosion outcome, and it shows that high-compactness progenitors above $\xi_{2.0}\sim 0.5$ lead to successful explosions both in multi-D, and 1D+ simulations, with the only exception of four progenitors simulated with GR1D+ and one progenitor simulated in 2D with FLASH. The former are a 25 $M_{\odot}$ progenitor at zero metallicity and the 23.2, 23.8, and 24 $M_{\odot}$ at a metallicity of $10^{-4}z_{\rm sol}$ from Woosley et al. (2002), where $z_{\rm sol}$ is solar metallicity. The latter is a 35 $M_{\odot}$ progenitor from Woosley & Heger (2007) simulated using a very stiff equation of state, i.e. the Skyrme-type EOS with effective nucleon mass $m^{*}$ = 0.55 from Eggenberger Andersen et al. (2021), and therefore it is not surprising that it does not explode.
Notice that if one were to use a stiffer equation of state, the overall curve $\dot{Q}_{\nu}^{\rm max}$ vs. $\xi_{2.0}$ would be shifted down, since stiff equations of state lead to colder neutrinospheres, and therefore lower neutrino heating (Eggenberger Andersen et al., 2021; Schneider et al., 2020). The few 2D FLASH simulations performed with stiffer equations of state (in particular, with lower effective nucleon masses) are compatible with this finding.
From Figure 3 we can confirm that modern multi-dimensional simulations clearly show that high-compactness progenitors explode. It is however important to define what "explosion" means in this case. All of the simulations analyzed in this paper only simulate up to 1-5 seconds after bounce, depending on the model. In all simulations of high-compactness progenitors (except for very stiff EOSs, as discussed above), the shock was successfully revived and expanded to radii of thousands of kilometers. Whether the shock will successfully break out of the star and, more importantly, what energy that explosion will be, is however not obvious. Several studies (Eggenberger Andersen et al., 2024; Sykes & Müller, 2024; Chan et al., 2018; Burrows et al., 2023; Powell et al., 2021; Burrows et al., 2024a) have shown that it is indeed possible to achieve an explosion (with energies between one-hundredth of a Bethe to a few Bethe) of these high-compactness progenitors. In particular, Sykes & Müller (2024) have shown that as long as a black hole forms after the shock has propagated past the sonic point within the envelope, the shock will break out of the star. With all of that considered, we can assume that simulations, where the shock has been successfully revived, yield an explosion. Therefore, we assume that all of the 2D and 3D simulations with compactness $\xi_{2.0}\gtrsim 0.5$ explode.
@@
It is also instructive to run two identical 1D+ models implemented in different codes (Couch et al., 2020; Boccioli et al., 2021). The 1D+ FLASH simulations of the KEPLER progenitors from Sukhbold et al. (2016) follow the same trend of the GR1D+ simulations, with a systematically slightly lower neutrino heating. The reason is that 1D+ models can be quite sensitive to the mixing-length parameter $\alpha_{\rm MLT}$. FLASH is a Newtonian code, whereas GR1D+ is fully relativistic. Therefore, the expression for the Brunt-Väisälä frequency is different, which results in different calibration of the mixing length parameter $\alpha_{\rm MLT}$ (Couch et al., 2020; Boccioli et al., 2021). The fact that the heating is consistently lower for all progenitors indicates that indeed if one were to use a slightly different value of $\alpha_{\rm MLT}$ (either in FLASH or GR1D+) one would get an even better agreement between the two 1D+ models.
@@
Figure 3: Maximum of the neutrino heating as a function of compactness. The bottom panel shows only 1D+, 2D, and 3D simulations, color-coded based on explosion outcome. All simulations (except for 3 edge cases) above $\xi_{2.0}\sim 0.5$ lead to successful explosions. In the top panel, we also include 1D simulations (in grey), and different symbols and colors represent different sets of simulations. All simulations adopt the SFHo EOS (except for the dark brown pluses(*), which indicate 2D FLASH simulations run with a few different EOSs). Typically, the stiffer the EOS the smaller the neutrino heating. The agreement between 1D+, 2D, and 3D simulations is remarkable. Three general trends can be seen in the Figure: the one followed by 1D simulations, shown in Eq. (11), the one followed by multi-D and 1D+ simulations for compactnesses below $\xi_{2.0}\sim 0.72$, shown in Eq. (12), and the even steeper one for compactnesses above $\xi_{2.0}\sim 0.72$, shown in Eq. (13). Different trends are illustrated by the two dashed lines and the dotted black line in the upper panel. Notice that these are not direct fits to $\dot{Q}_{\nu}^{\rm max}$ vs. $\xi_{2.0}$, and the details of how they were derived are given in Appendix B.

### v2

@@ 2.2 Neutrino heating in 1D+ models
The main "extra" 1D+ contribution to the spherically symmetric Euler equations is $\rho\epsilon_{g}-\nabla\cdot F_{e}-\nabla\cdot F_{K}$, as can be seen from Eq. (7). To quantify how much neutrino-driven convection changes the overall heating in the gain region, we selected 9 different 1D and 1D+ simulations from Boccioli & Fragione (2024) for 9 different values of compactness $\xi_{2.0}$. The 1D+ and 1D simulations only differ for the inclusion of the STIR turbulent model, with everything else being equal. The properties of the selected progenitors are shown in Table 1. These simulations were run with GR1D (O’Connor & Ott, 2010), an open-source spherically symmetric, general relativistic code with an M1 neutrino transport and neutrino opacities from the open-source library NuLib, described in O’Connor (2015). The publicly available code [^1] https://github.com/evanoconnor/GR1D has been modified with the addition of STIR (Couch et al., 2020; Boccioli et al., 2021) described in the previous section. More details on the numerical setup adopted for GR1D can be found in Boccioli et al. (2021) and Boccioli & Fragione (2024).
To illustrate the effects of convection and neutrino heating in the STIR model we show in Figure 1 the time and radial dependence of several key quantities for these 9 1D+ and 1D simulations. Notice that, to better show how much each quantity contributes to the overall heating, each quantity is multiplied by the volume of each zone. The first two rows show the (negative) divergence of the internal and turbulent energy convective fluxes (i.e. $-\nabla\cdot F_{e}$ and $-\nabla\cdot F_{K}$) in the 1D+ simulations. As mentioned at the start of this section, these can be thought of as additional source terms in Eq. (7). The term $-\nabla\cdot F_{e}$ represents the net change in internal energy resulting from the transport of hot bubbles closer to the shock. One would expect that below the gain region, i.e. where the entropy gradient vanishes, $-\nabla\cdot F_{e}$ would vanish. However, this is not the case and it has been pointed out by Müller (2019) that this is indeed a potential issue of the STIR model. This relatively small inconsistency is caused by an inadequate expression for the convective flux $F_{e}$, which instead of being expressed in term of $\nabla\epsilon$, as done in STIR, should instead depend on $\nabla\epsilon+P\nabla(1/\rho)$.
In the context of this paper, we allow for this inconsistency that, as can be inferred by the upper panels of Figure 1, causes an overall positive net heating when $-\nabla\cdot F_{e}$ is integrated over the gain region. Importantly, even if $-\nabla\cdot F_{e}$ would vanish at the bottom of the gain region, causing vanishing net heating within the gain region, the material behind the shock would still experience positive net heating, which is what is most important for shock revival. As shown in the second row, $-\nabla\cdot F_{K}$ is instead negative throughout the gain region, which indicates that turbulent energy is taken from the center of the region (i.e. the minimum of $-\nabla\cdot F_{K}$, which corresponds to the maximum of $v_{\rm turb}$) and transported away from it. Turbulent energy is therefore transported toward the shock or toward the bottom of the gain region, depending on whether it is located above or below the radial coordinate of the maximum of $v_{\rm turb}$ (i.e. the minimum of $-\nabla\cdot F_{K}$). At the edges of the convective region $-\nabla\cdot F_{K}$ is instead positive, since $v_{\rm turb}$ vanishes. As expected, $-\nabla\cdot F_{K}$ is roughly zero when integrated over the gain region.
The third row of Figure 1 shows $\rho\epsilon_{g}$, which has a similar radial dependence of the neutrino heating shown on the fourth row. This is consistent with the fact that neutrino heating determines the entropy profile in the gain region, which in turn determines the strength of convection. So far, the quantities in the first four rows were discussed for the 1D+ simulations. The fifth row shows instead the neutrino heating for the respective 1D simulations. As one can see, the radial dependence and magnitude of $\rho\epsilon_{\nu}$ are the same in 1D and 1D+ models. However, as we will show in detail in Section 4 the crucial difference is the shock radius and hence the size of the gain region, which is much larger in 1D+ models. Therefore, the total net heating in the gain region is much larger in the 1D+ simulations because the size of the gain region is larger. We will come back to this later in the manuscript.
One exception is the last snapshot at $\sim 220$ ms, which shows a different radial dependence in the 1D simulation, where $\rho\epsilon_{\nu}\,dV$ quickly increases behind the shock. When the shock recedes, as is the case for the last snapshot, the dependence of the net neutrino heating becomes inversely proportional to radius. Since $dV\propto r^{2}$, overall, one can see a linear increase in $\rho\dot{\epsilon}_{\nu}dV$. The higher compactness progenitors will have the same linear dependence on radius, although steeper, since luminosities and neutrino energies are larger. The reason why the net neutrino heating is inversely proportional to radius can be understood by the fact that, in general, neutrino heating is proportional to $1/r^{2}$, whereas neutrino cooling is proportional to $1/r^{6}$ in a radiation-dominated region. When the shock recedes, the entire gain region will be located at small radii, where the cooling is larger, hence softening the overall dependence of net neutrino heating from $1/r^{2}$ to $\sim 1/r$.
@@
Table 1: Summary of the properties of the progenitors used in Section 4. The simulations were performed in Boccioli & Fragione (2024).
Progenitors with higher compactness are characterized by shallower density profiles, and therefore larger mass accretion rates. As a consequence, they generate larger neutrino heating, as can be seen in Figure 1, while at the same time, they also generate larger ram pressure onto the shock. In the remainder of this paper, we will show that high-compactness progenitors, in the presence of $\nu$-driven convection, are able to generate large neutrino heating very efficiently, which has a direct impact on explodability.
@@
which represent the rates of turbulent generation and dissipation, respectively. Notice that, in principle, one expects $E_{g}\approx E_{k}$, since the overall turbulent energy generated should balance the turbulent energy dissipated, as shown by Murphy et al. (2013). This is however not necessarily true in STIR, and instead, $E_{g}$ is about $10\%$ larger than $E_{k}$. The origin of this difference lies in the fact that when one integrates Eq. (2) over the gain region, $\nabla\cdot F_{e}$ is non-zero as mentioned above. Instead, the first and second terms of Eq. (2) are both roughly zero when integrated over the gain region, a sign that during the accretion phase, the gain region is in a quasi-steady-state.
It should be highlighted that $\epsilon_{g}$ and $\epsilon_{k}$ are rates of turbulent dissipation and generation, and therefore $\epsilon_{g,k}$ and $E_{g,k}$ are dimensionally equivalent to $\dot{\epsilon}_{\nu}$ and $\dot{Q}_{\nu}$, respectively. This makes the notation slightly ambiguous since one might expect those quantities to be defined as $\dot{\epsilon}_{g,k}$ and $\dot{E}_{g,k}$. However, to be consistent with previous literature (Murphy et al., 2013; Mabanta & Murphy, 2018; Couch et al., 2020), we decided to omit the dot.
@@ 3 Comparison between STIR and multi-dimensional simulations
To illustrate how the STIR model compares to multi-dimensional simulations, we compare neutrino heating in STIR with neutrino heating in state-of-the-art 2D and 3D Fornax and 2D FLASH simulations. Specifically, we will analyze 98 2D and 20 3D Fornax simulations and 55 2D FLASH simulations. The 2D Fornax simulations (Vartanyan & Burrows, 2023) were performed on solar metallicity progenitors from Sukhbold et al. (2016) with initial masses between 9 and 12 $M_{\odot}$, and on solar metallicity progenitors from Sukhbold et al. (2018) with initial masses between 12 and 26.99 $M_{\odot}$. The 3D Fornax simulations (Burrows et al., 2020) were performed on solar metallicity progenitors from Sukhbold et al. (2016), with masses ranging from 12 to 60 $M_{\odot}$, as well as a $25M_{\odot}$ progenitor from Sukhbold et al. (2018). All of the Fornax simulations adopted the SFHo equation of state. The 2D FLASH simulations were performed on different progenitors with masses from $12-60M_{\odot}$ from Sukhbold et al. (2016) as well as a few from Woosley & Heger (2007) and some with artificially reduced mass loss from Sukhbold et al. (2018). Of these progenitors, 37 were simulated using the SFHo EOS, whereas the rest were simulated using other EOSs from Schneider et al. (2019). The FLASH simulations are collected from various studies but also from publications in preparation (Eggenberger Andersen et al., 2021, 2025; Li, 2024; Andresen et al., 2024).
@@
Several groups have performed comparisons between different codes, although usually it has been done for simulations in either 1D (O’Connor et al., 2018), 2D (Pan et al., 2019), or 3D (Cabezón et al., 2018; Glas et al., 2019), and more rarely have 1D simulations been thoroughly compared to multi-dimensional ones (Richers et al., 2017; Just et al., 2018; Nagakura et al., 2020). The latter is a particularly difficult comparison to perform given that 1D simulations do not explode, whereas multi-dimensional ones often do, and therefore the most significant differences between 1D and multi-D results can be attributed to the different outcome rather than specific multi-dimensional effects.
@@
In this section, we focus on comparing neutrino heating in 1D, 1D+, 2D, and 3D simulations. We adopt simulations already performed with other studies in mind, which means that neutrino transport, opacities, and overall simulation setups are not necessarily the same across codes. These differences prevent a detailed and thorough comparison study, which is however beyond the scope of this paper. Therefore, some of the discrepancies between 1D+ and multi-D, as well as discrepancies between the 2D FLASH and Fornax simulations, are to be expected. The largest differences are most likely due to neutrino opacities and transport methods, and we refer to the original papers listed at the beginning of this section for details regarding the specific neutrino treatments in the different codes.
@@
In general, neutrino heating (but also several other quantities) in 2D simulations tends to be a lot noisier than in 3D, even though the neutrino luminosities of 2D and 3D Fornax simulations agree very well at early times $\lesssim 1$s (Vartanyan & Burrows, 2023). Because of the imposed axial symmetry, 2D simulations are characterized by much larger structures than their 3D counterparts. The average in the phi direction in 3D simulations is what smooths out the large oscillations. Indeed, in a 3D simulation, the neutrino heating calculated in different wedges at different values of phi, show large variations in neutrino heating up to $\sim 20\%$, which are then smoothed out by the phi-average. This is the main reason why neutrino heating in 2D simulations does not average out as much compared to 3D.
@@
For the low-compactness progenitor, the 2D Fornax simulations produce a larger neutrino heating at early times, compared to both the 2D FLASH and 3D Fornax simulations, which instead seem to agree relatively well. This is seen for all of the lower-compactness progenitors in the 2D Fornax models compared to 3D Fornax and 2D FLASH models (see Section 3 for more details). The larger neutrino heating in the 2D Fornax simulations can be caused by a number of factors, such as the different geometry, grid resolution, and neutrino opacities and transport schemes. Here we simply describe these differences and argue what their potential impact might be, but a thorough analysis of this discrepancy is beyond the scope of this work.
FLASH and Fornax are both Newtonian codes, that include general relativistic corrections to the monopole gravitational potential using Case A of Marek et al. (2006). The grid resolutions for all the multi-D simulations are overall quite similar, with the smallest radial zone being $\sim 500$ m. The hydrodynamics solver is however slightly different. FLASH simulations adopt an unsplit hydrodynamic solver, Spark (Couch et al., 2021), which makes use of an adaptive mesh refinement grid and uses fifth order WENO reconstruction. Fornax instead is a finite-volume code (Skinner et al., 2019) adopting a “dendritic grid", i.e. a static-mesh refinement near the center and the poles, and adopts a custom reconstruction algorithm described in Burrows et al. (2020). Both codes use an HLLC Riemann solver that reduces to HLLE near shocks. The neutrino transport scheme is also very similar, and both codes adopt a multi-dimensional, multi-energy transport algorithm that solves the Boltzmann equations using an M1-scheme and the maximum-entropy Minerbo closure (Minerbo, 1978). Fornax solves the transport equations in the co-moving frame, whereas FLASH solves them in the laboratory frame. The opacities adopted are, for the most part, the same, but we refer to the method papers cited above for more details. The most significant difference between 2D Fornax models and 2D FLASH models is perhaps the adoption in Fornax of a Kompaneets approach for inelastic neutrino-nucleon scattering (Wang & Burrows, 2020), whereas the FLASH simulations only consider the elastic version of this process. This is also the main difference (other than, of course, geometry) between the 2D and 3D Fornax simulations. That said, it seems however unlikely that this could cause such large differences in neutrino heating at early times, as seen for the 12.0 and 12.15 $M_{\odot}$ progenitors, although more comparison work in this direction is needed. Notably, no thorough comparisons among state-of-the-art multi-dimensional codes regarding the impact of different neutrino opacities have been carried out yet, due to their complexity.
@@
An important detail worth mentioning is that, depending on how the gain region is defined, certain computational zones might be included or excluded from the integral in Eq. (8). Therefore, it is important to define a cut in density (i.e., exclude zones with densities above $3\times 10^{10}\ {\rm g/cm^{3}}$) to exclude zones inside the PNS which might experience large numerical fluctuations in net heating. However, this is not enough since some zones just above the PNS might still be included in the integral, even though they do not influence the neutrino heating in the vicinity of the shock, which ultimately is what Eq. (8) should reflect. Requiring the specific entropy per baryon to be above 6 ensures that only material in the shocked region is not included in the integral. Not including this extra condition was seen to increase the neutrino heating in the 2D Fornax simulations by up to $\sim 10\%$, and it is therefore important to take it into account.
@@ 3.1 Compactness dependence of neutrino heating
In this section, we will focus on several suites of simulations in 1D, 1D+, 2D and 3D, to study how neutrino heating depends on the progenitor compactness. For that purpose, we will use all of the 2D and 3D simulations described above. In addition to that, we selected several suites of both 1D and 1D+ simulations ran with GR1D: 309 FRANEC progenitors (Chieffi et al., 2021) from Boccioli et al. (2023); 341 KEPLER progenitors (Woosley et al., 2002; Sukhbold et al., 2016) from Boccioli & Fragione (2024); 88 (only 1D+ simulations) KEPLER progenitors (Sukhbold et al., 2018), which are the same progenitors as the 2D Fornax simulations by Vartanyan & Burrows (2023). Finally, we also selected 148 1D FLASH simulations (Segerlund et al., 2021) of KEPLER progenitors from Sukhbold et al. (2016) with masses between 9 and 120 $M_{\odot}$, as well as 40 1D+ FLASH simulations (Couch et al., 2020) with masses from 9 to $30M_{\odot}$. In total, we therefore have 173 multi-D simulations, 778 1D+ simulations, and 798 1D simulations, for a total of $\sim 1500$ simulations.
@@
Figure 3 shows $\dot{Q}_{\nu}^{\rm max}$ as a function of the pre-supernova compactness for the $\sim 1500$ simulations described above. The bottom panel is color-coded based on explosion outcome, and it shows that high-compactness progenitors above $\xi_{2.0}\sim 0.5$ lead to successful explosions both in multi-D, and 1D+ simulations, with the only exception of four progenitors simulated with GR1D+ and one progenitor simulated in 2D with FLASH. The former are a 25 $M_{\odot}$ progenitor at zero metallicity and the 23.2, 23.8, and 24 $M_{\odot}$ at a metallicity of $10^{-4}z_{\odot}$ from Woosley et al. (2002), where $z_{\odot}$ is solar metallicity. The latter is a 35 $M_{\odot}$ progenitor from Woosley & Heger (2007) simulated using a very stiff equation of state, i.e. the Skyrme-type EOS with effective nucleon mass $m^{*}$ = 0.55 from Eggenberger Andersen et al. (2021), and therefore it is not surprising that it does not explode.
Notice that if one were to use a stiffer equation of state, the overall curve $\dot{Q}_{\nu}^{\rm max}$ vs. $\xi_{2.0}$ would be shifted down, since stiff equations of state lead to colder neutrinospheres, and therefore lower neutrino heating (Schneider et al., 2020; Eggenberger Andersen et al., 2021). The few 2D FLASH simulations performed with stiffer equations of state (in particular, with lower effective nucleon masses) are compatible with this finding.
From Figure 3 we can confirm that modern multi-dimensional simulations clearly show that high-compactness progenitors explode. It is however important to define what "explosion" means in this case. All of the simulations analyzed in this paper only simulate up to 1-5 seconds after bounce, depending on the model. In all simulations of high-compactness progenitors (except for very stiff EOSs, as discussed above), the shock was successfully revived and expanded to radii of thousands of kilometers. Whether the shock will successfully break out of the star and, more importantly, what energy that explosion will be, is however not obvious. Several studies (Chan et al., 2018; Powell et al., 2021; Burrows et al., 2023; Eggenberger Andersen et al., 2025; Sykes & Müller, 2025; Burrows et al., 2024a) have shown that it is indeed possible to achieve an explosion (with energies between one-hundredth of a Bethe to a few Bethe) of these high-compactness progenitors. In particular, Sykes & Müller (2025) have shown that as long as a black hole forms after the shock has propagated past the sonic point within the envelope, the shock will break out of the star. With all of that considered, we can assume that simulations, where the shock has been successfully revived, yield an explosion. Therefore, we assume that all of the 2D and 3D simulations with compactness $\xi_{2.0}\gtrsim 0.5$ explode.
@@
It is also instructive to run two identical 1D+ models implemented in different codes (Couch et al., 2020; Boccioli et al., 2021). The 1D+ FLASH simulations of the KEPLER progenitors from Sukhbold et al. (2016) follow the same trend of the GR1D+ simulations, with a systematically slightly lower neutrino heating. The reason is that 1D+ models can be quite sensitive to the mixing-length parameter $\alpha_{\rm MLT}$. Moreover, FLASH adopts modified Newtonian gravity, whereas GR1D+ is a general relativistic code. Therefore, the expression for the Brunt-Väisälä frequency is different, which results in different calibration of the mixing length parameter $\alpha_{\rm MLT}$, which is $\sim 1.23-1.27$ in FLASH (Couch et al., 2020) and $\sim 1.48-1.52$ in GR1D (Boccioli et al., 2021). The fact that the heating is consistently lower in FLASH for all progenitors indicates that if one were to use a slightly higher value of $\alpha_{\rm MLT}$ in FLASH, or slightly lower in GR1D+, within the ranges mentioned above, one would get an even better agreement between the two 1D+ models.
@@
Figure 3: Maximum of the neutrino heating as a function of compactness. The bottom panel shows only 1D+, 2D, and 3D simulations, color-coded based on explosion outcome. Notice that the failed explosions have been plotted on top of the successful ones, and since several points overlap, it looks like the number of failed explosions is much higher than it is in reality. All simulations (except for four 1D+ and one 2D simulation discussed in the text)
above $\xi_{2.0}\sim 0.5$ lead to successful explosions. In the top panel, we also include 1D simulations (in grey), and different symbols and colors represent different sets of simulations. All simulations adopt the SFHo EOS (except for the dark brown pluses(*), which indicate 2D FLASH simulations run with a few different EOSs). Typically, the stiffer the EOS the smaller the neutrino heating. The agreement between 1D+, 2D, and 3D simulations is remarkable. Three general trends can be seen in the Figure: the one followed by 1D simulations, shown in Eq. (12), the one followed by multi-D and 1D+ simulations for compactnesses below $\xi_{2.0}\sim 0.72$, shown in Eq. (13), and the even steeper one for compactnesses above $\xi_{2.0}\sim 0.72$, shown in Eq. (14). Different trends are illustrated by the two dashed lines and the dotted black line in the upper panel. Notice that these are not direct fits to $\dot{Q}_{\nu}^{\rm max}$ vs. $\xi_{2.0}$, and the details of how they were derived are given in Appendix B.